OpenStudy (anonymous):

In a right triangle ABC, the legs are AB = 5 and BC = 12, and in the right triangle DEF, the hypotenuse is DE = 10, and EF = 8. Which is DF + AC? Answer 19 15 13 12

OpenStudy (anonymous):

In a right angle triangle Hypotenuse^2=Base^2+Altitude^2 AC^2=5^2+12^2=169=13^2 AC=13 DE^2=EF^2+DF^2 DF^2=DE^2-EF^2=100-64=36 DF=6 AC+DF=13+6=19